# Inflation, the Hubble Tension and Early Dark Energy: an alternative overview

### v1

@@ I Introduction
The significant advancements in observational astrophysics and cosmology, culminating in a vast array of experiments probing different cosmic epochs and scales Bennett et al. (2013); Hinshaw et al. (2013); Aghanim et al. (2020a, b); Ruhl et al. (2004); Dutcher et al. (2021); Choi et al. (2020); Aiola et al. (2020); Madhavacheril et al. (2023); Dawson et al. (2013); Anderson et al. (2014); Delubac et al. (2015); Alam et al. (2017); Beutler et al. (2014); Alam et al. (2021); Tegmark et al. (2004); Seljak et al. (2005); Tegmark et al. (2006); Betoule et al. (2014); Abbott et al. (2016); Troxel et al. (2018); Abbott et al. (2022, 2023a); Asgari et al. (2021); de Jong et al. (2013); Tröster et al. (2021); Abbott et al. (2023b); Scolnic et al. (2018); Brout et al. (2022); Abbott et al. (2017a, 2019, b), have allowed for precise constraints on the inflationary Universe and its related observables. Over the last few decades, several inflationary models and theories have been tested against a wide range of available data, including Cosmic Microwave Background (CMB), Big Bang Nucleosynthesis (BBN), and Gravitational Wave (GW) measurements [^1] With no claim to be exhaustive, for works in this direction see, e.g., Refs. Leach et al. (2002); Boubekeur and Lyth (2005); Martin and Ringeval (2006); Moss and Graham (2007); Bezrukov et al. (2011); Zhao and Huang (2011); Martin et al. (2014a, b, c); Carrillo-González et al. (2014); Creminelli et al. (2014); Di Valentino and Mersini-Houghton (2017a, b); Campista et al. (2017); Giarè et al. (2019); Forconi et al. (2021); Dai and Zhu (2020); Baumann et al. (2016); Odintsov et al. (2021); Giarè et al. (2021a); Oikonomou (2021); Odintsov et al. (2022); Namba et al. (2016); Peloso et al. (2016); Pi et al. (2019); Özsoy (2021); Stewart and Brandenberger (2008); Mukohyama et al. (2014); Giovannini (2016, 2018a, 2019, 2018b); Giarè and Melchiorri (2021); Giarè and Renzi (2020); Giarè et al. (2021a, 2023a); Baumgart et al. (2022); Franciolini et al. (2019); D’Eramo and Schmitz (2019); Giarè et al. (2019); Caldwell et al. (2019); Clarke et al. (2020); Caprini and Figueroa (2018); Allen and Romano (1999); Smith et al. (2006); Boyle and Buonanno (2008); Kuroyanagi et al. (2015); Ben-Dayan et al. (2019); Aich et al. (2020); Cabass et al. (2016); Vagnozzi (2021); Benetti et al. (2022); Calcagni and Kuroyanagi (2021); Oikonomou (2023a); Barrow et al. (1993); Peng et al. (2021); Ota et al. (2023); Odintsov and Oikonomou (2022a); Baumgart et al. (2022); Capurri et al. (2020); Cañas Herrera and Renzi (2021); Odintsov et al. (2023); Oikonomou (2023b); Fronimos and Venikoudis (2023, 2023); Cai (2023); Oikonomou (2022); Gangopadhyay et al. (2023); Odintsov and Oikonomou (2022b); Odintsov et al. (2022, 2020); Galloni et al. (2023); Braglia et al. (2023); Giarè et al. (2023b); Antoniadis et al. (2023); Agazie et al. (2023); Afzal et al. (2023); Vagnozzi (2023a); Oikonomou (2023c); Iacconi et al. (2020, 2022, 2023); Santos et al. (2023); Pozo et al. (2023); Giarè et al. (2024a); Cecchini et al. (2024) and references therein. See also the recent Ref. Martin et al. (2024) for a detailed Bayesian model comparison involving nearly three hundred models of single-field slow-roll inflation. . Despite embedding inflation within a more fundamental theory remains an open problem Lyth and Riotto (1999); Linde (2008); Martin et al. (2014a); Baumann and McAllister (2015) and a large plethora of proposed models and mechanisms can be deemed equally viable in describing current observations Martin et al. (2014a); Akrami et al. (2020a), a few general guidelines can be drawn from existing data.
@@
Firstly, the aforementioned discrepancies among different CMB experiments primarily involve inflationary parameters Forconi et al. (2021); Giarè et al. (2023d); Di Valentino et al. (2022); Gariazzo et al. (2024). Unexpectedly, small-scale CMB measurements provided by the Data Release 4 of the Atacama Cosmology Telescope (ACT) show an agreement with a scale invariant Harrison-Zel’dovich spectrum ($n_{s}=1.009\pm 0.015$), introducing a tension with a significance of $99.3\%$ CL with the results from the Planck satellite Giarè et al. (2023d). If we set aside observational systematics and consider the differences between ACT and Planck as genuine, the analysis of small-scale CMB data lead to different predictions for inflationary models Giarè et al. (2023c).
@@
As many have argued, resolving the Hubble tension might require a paradigm shift in cosmology, introducing new physics into the standard cosmological model. Although it seems unlikely that the new physics needed to resolve the Hubble tension would come entirely from the inflationary sector of the theory, in general, considering new physics within the cosmological model could lead to changes in the values of cosmological parameters inferred from data. In this regard, I would like to highlight one more time that, just like with $H_{0}$, the results concerning inflationary parameters discussed so far (most typically $n_{s}$) are derived from the analysis of CMB data within the standard cosmological model. Considering different models may affect – although indirectly – our understanding of the inflationary Universe. A striking example supporting this claim comes from Early Dark Energy (EDE), a highly studied theoretical extension representing a possible solution to the Hubble tension Wetterich (2004); Doran and Robbers (2006); Hollenstein et al. (2009); Calabrese et al. (2011a, b, b); Pettorino et al. (2013); Archidiacono et al. (2014); Poulin et al. (2023, 2018a, 2018b); Smith et al. (2020); Niedermann and Sloth (2021, 2020); Murgia et al. (2021); Ye and Piao (2020); Klypin et al. (2021); Hill et al. (2020); Herold et al. (2022); Herold and Ferreira (2023); Reeves et al. (2023); Jiang and Piao (2022); Simon et al. (2023); Smith et al. (2022); Kamionkowski and Riess (2023); Niedermann and Sloth (2023); Cruz et al. (2023); Eskilt et al. (2023); Smith and Poulin (2023); Sharma et al. (2023); Efstathiou et al. (2023); Gsponer et al. (2023); Goldstein et al. (2023); Forconi et al. (2023). No matter whether analyzed in light of Planck or ACT data, this model consistently yields a higher fitting value for the scalar spectral index Di Valentino et al. (2018); Ye et al. (2021, 2022); Jiang and Piao (2022); Jiang et al. (2022); Takahashi and Yin (2022); Jiang et al. (2023); Peng and Piao (2023); Forconi et al. (2023). This suggests that, in EDE cosmology, resolving the Hubble tension requires moving towards the same direction favored by small-scale CMB experiments (i.e., towards larger values of $n_{s}$) [^5] Notably, evidence at a 3 standard deviation level in favor of EDE is obtained precisely from such experiments, reflecting, among other things, this preference for a higher $n_{s}$ Hill et al. (2020)..
@@ II.1 Early-time new physics
where $N_{\rm eff}$ is the effective number of relativistic particles, while $\Omega_{c}$, $\Omega_{b}$, and $\Omega_{\gamma}$ are the cold dark matter, baryons, and photons energy densities, respectively. The energy-density of various species can be accurately measured within a narrow margin of error, both through the effects in the CMB and through independent observations such as the BBN (for $\Omega_{b}$) or BAO measurements (for $\Omega_{c}$). Therefore, as also highlighted in Ref. Vagnozzi (2020), a simple (yet physically motivated) method to parameterize new physics able to increase the rate of expansion of the Universe before recombination consists of increasing the value of the effective number of relativistic particles, $N_{\text{eff}}$. Within the standard cosmological model, this parameter takes the reference value of $N_{\rm eff}=3.044$ Mangano et al. (2005); Akita and Yamaguchi (2020); Froustey et al. (2020); Bennett et al. (2021), accounting for three different families of relativistic neutrinos along with an additional contribution arising from non-instantaneous neutrino decoupling. However, given its flexibility in quantifying each relativistic component in the early Universe, this value can be changed by a wide range of phenomenological extensions spanning from modifications to the standard model of particle physics that involve new relativistic degrees of freedom active in the early Universe, to models that incorporate additional scalar degrees of freedom or significant graviton cosmic background, see e.g., Refs Giarè et al. (2021b); D’Eramo et al. (2022); Giarè et al. (2022, 2023a).
@@ II.2 Late-time new physics
where $\Omega_{m}=\Omega_{b}+\Omega_{c}+\Omega_{\nu}$ represents the matter energy-density (encompassing baryons, cold dark matter, and neutrinos), $\Omega_{k}$ denotes the curvature density parameter, and $\Omega_{\rm{DE}}$ describes the energy-density component associated with Dark Energy (DE). Notice that I assume the DE equation of state $w\equiv P_{\rm{DE}}/\rho_{\rm{DE}}$ (i.e., the ratio between the pressure $P_{\text{DE}}$ and energy-density $\rho_{\rm{DE}}$ of DE) to be constant. Within the standard $\Lambda$CDM model of cosmology, this parameter is fixed to $w=-1$, corresponding to a cosmological constant case $\Omega_{\rm{DE}}=\Omega_{\Lambda}$. Within $\Lambda$CDM the spatial background geometry is assumed to be flat, $\Omega_{k}=0$. Therefore, if we aim to use the simple Eq. (5) to parameterize the effects of new physics beyond the standard cosmological model, we are left with only two possibilities: working on the background geometry or on the DE equation of state [^7] Notice that many other attempts to solve the Hubble tension in the context of late-time new physics have been discussed in the literature, introducing modifications to the background dynamics parameterized by Eq. (5). Just to mention a few concrete examples, a non-exhaustive list of possibilities involves considering modified gravity theories and or exotic/phenomenological theoretical frameworks such as Interacting Dark Energy or sign-switching cosmology. With no claim of covering all possibilities, for works in this direction, the interested reader can refer to, e.g., Refs. Gavela et al. (2009); Di Valentino et al. (2017); Kumar and Nunes (2017); Wang et al. (2016); Di Valentino et al. (2020a); Pan et al. (2019a); Yang et al. (2018); Murgia et al. (2016); Yang et al. (2019); Pan et al. (2019b); Di Valentino et al. (2020b); Lucca and Hooper (2020); Gómez-Valent et al. (2020); Di Valentino (2021); Yang et al. (2021a); Gariazzo et al. (2022); Bernui et al. (2023); Zhai et al. (2023); Giarè et al. (2024b); Montani et al. (2023a); Schiavone et al. (2023); Montani et al. (2023b); Akarsu et al. (2021, 2023a, 2023b); Gómez-Valent et al. (2024) and references therein.. Both of these parameters can be well constrained by a multitude of observational data and have been the subject of intense attention and debate for a variety of reasons [^8] See, e.g., Refs. Park and Ratra (2019); Handley (2021); Di Valentino et al. (2019); Efstathiou and Gratton (2020); Di Valentino et al. (2021c); Benisty and Staicova (2021); Vagnozzi et al. (2021a, b); Di Valentino et al. (2021d); Yang et al. (2021a); Cao et al. (2021); Dhawan et al. (2021); Dinda (2022); Gonzalez et al. (2021); Akarsu et al. (2023c); Cao and Ratra (2022); Glanville et al. (2022); Bel et al. (2022); Yang et al. (2023); Stevens et al. (2023); Favale et al. (2023) for discussions surrounding the curvature parameter and Refs. Escamilla et al. (2023); Aghanim et al. (2020b); Alam et al. (2021); Semenaite et al. (2023); Carrilho et al. (2023); Chudaykin et al. (2021); D’Amico et al. (2021); Brieden et al. (2022); Scolnic et al. (2018); Brout et al. (2022); Abbott et al. (2019); Tröster et al. (2021); Abbott et al. (2023a); Moresco et al. (2016); Vagnozzi et al. (2021b); Yang et al. (2021b); Di Valentino (2021); Moresco et al. (2022); Vagnozzi et al. (2022); Bargiacchi et al. (2022); Moresco et al. (2022); Grillo et al. (2020); Cao et al. (2022); Hogg (2023); Adame et al. (2024) for recent constraints on the DE equation of state from a multitude of astrophysical and cosmological probes.. Again, I will not take part in the debate as my goal is to parameterize new physics able to alleviate the Hubble tension and study its implications for inflation. Since one of the most typical predictions of inflation is that the observable Universe should be locally flat, I fix the background geometry to $\Omega_{k}=0$, focusing on the potential realizations of new theories capable of predicting non-standard values for $w$.

### v2

@@ I Introduction
The significant advancements in observational astrophysics and cosmology, culminating in a vast array of experiments probing different cosmic epochs and scales Bennett et al. (2013); Hinshaw et al. (2013); Aghanim et al. (2020a, b); Ruhl et al. (2004); Dutcher et al. (2021); Choi et al. (2020); Aiola et al. (2020); Madhavacheril et al. (2023); Dawson et al. (2013); Anderson et al. (2014); Delubac et al. (2015); Alam et al. (2017); Beutler et al. (2014); Alam et al. (2021); Tegmark et al. (2004); Seljak et al. (2005); Tegmark et al. (2006); Betoule et al. (2014); Abbott et al. (2016); Troxel et al. (2018); Abbott et al. (2022, 2023a); Asgari et al. (2021); de Jong et al. (2013); Tröster et al. (2021); Abbott et al. (2023b); Scolnic et al. (2018); Brout et al. (2022); Abbott et al. (2017a, 2019, b), have allowed for precise constraints on the inflationary Universe and its related observables. Over the last few decades, several inflationary models and theories have been tested against a wide range of available data, including Cosmic Microwave Background (CMB), Big Bang Nucleosynthesis (BBN), and Gravitational Wave (GW) measurements [^1] With no claim to be exhaustive, for works in this direction see, e.g., Refs. Leach et al. (2002); Boubekeur and Lyth (2005); Martin and Ringeval (2006); Moss and Graham (2007); Bezrukov et al. (2011); Zhao and Huang (2011); Martin et al. (2014a, b, c); Carrillo-González et al. (2014); Creminelli et al. (2014); Di Valentino and Mersini-Houghton (2017a, b); Campista et al. (2017); Giarè et al. (2019); Forconi et al. (2021); Dai and Zhu (2020); Baumann et al. (2016); Odintsov et al. (2021); Giarè et al. (2021a); Oikonomou (2021); Odintsov et al. (2022); Namba et al. (2016); Peloso et al. (2016); Pi et al. (2019); Özsoy (2021); Stewart and Brandenberger (2008); Mukohyama et al. (2014); Giovannini (2016, 2018a, 2019, 2018b); Giarè and Melchiorri (2021); Giarè and Renzi (2020); Giarè et al. (2021a, 2023a); Baumgart et al. (2022); Franciolini et al. (2019); D’Eramo and Schmitz (2019); Giarè et al. (2019); Caldwell et al. (2019); Clarke et al. (2020); Caprini and Figueroa (2018); Allen and Romano (1999); Smith et al. (2006); Boyle and Buonanno (2008); Kuroyanagi et al. (2015); Ben-Dayan et al. (2019); Aich et al. (2020); Cabass et al. (2016); Vagnozzi (2021); Benetti et al. (2022); Calcagni and Kuroyanagi (2021); Oikonomou (2023a); Barrow et al. (1993); Peng et al. (2021); Ota et al. (2023); Odintsov and Oikonomou (2022a); Baumgart et al. (2022); Capurri et al. (2020); Cañas Herrera and Renzi (2021); Odintsov et al. (2023); Oikonomou (2023b); Fronimos and Venikoudis (2023, 2023); Cai (2023); Oikonomou (2022); Gangopadhyay et al. (2023); Odintsov and Oikonomou (2022b); Odintsov et al. (2022, 2020); Galloni et al. (2023); Braglia et al. (2023); Giarè et al. (2023b); Antoniadis et al. (2023); Agazie et al. (2023); Afzal et al. (2023); Vagnozzi (2023a); Oikonomou (2023c); Iacconi et al. (2020, 2022, 2023); Santos et al. (2023); Pozo et al. (2023); Giarè et al. (2024a); Cecchini et al. (2024); Wang et al. (2024) and references therein. See also the recent Ref. Martin et al. (2024) for a detailed Bayesian model comparison involving nearly three hundred models of single-field slow-roll inflation. . Despite embedding inflation within a more fundamental theory remains an open problem Lyth and Riotto (1999); Linde (2008); Martin et al. (2014a); Baumann and McAllister (2015) and a large plethora of proposed models and mechanisms can be deemed equally viable in describing current observations Martin et al. (2014a); Akrami et al. (2020a), a few general guidelines can be drawn from existing data.
@@
Firstly, the aforementioned discrepancies among different CMB experiments primarily involve inflationary parameters Forconi et al. (2021); Giarè et al. (2023d); Di Valentino et al. (2022); Gariazzo et al. (2024). Unexpectedly, small-scale CMB measurements provided by the Data Release 4 of the Atacama Cosmology Telescope (ACT) show an agreement with a scale invariant Harrison-Zel’dovich spectrum ($n_{s}=1.009\pm 0.015$), introducing a tension with a significance of $99.3\%$ confidence level with the results from the Planck satellite Giarè et al. (2023d). If we set aside observational systematics and consider the differences between ACT and Planck as genuine, the analysis of small-scale CMB data lead to different predictions for inflationary models Giarè et al. (2023c).
@@
As many have argued, resolving the Hubble tension might require a paradigm shift in cosmology, introducing new physics into the standard cosmological model. Although it seems unlikely that the new physics needed to resolve the Hubble tension would come entirely from the inflationary sector of the theory, in general, considering new physics within the cosmological model could lead to changes in the values of cosmological parameters inferred from data. In this regard, I would like to highlight one more time that, just like with $H_{0}$, the results concerning inflationary parameters discussed so far (most typically $n_{s}$) are derived from the analysis of CMB data within the standard cosmological model. Considering different models may affect – although indirectly – our understanding of the inflationary Universe. A striking example supporting this claim comes from Early Dark Energy (EDE), a highly studied theoretical extension representing a possible solution to the Hubble tension Wetterich (2004); Doran and Robbers (2006); Hollenstein et al. (2009); Calabrese et al. (2011a, b, b); Pettorino et al. (2013); Archidiacono et al. (2014); Poulin et al. (2023, 2018a, 2018b); Smith et al. (2020); Niedermann and Sloth (2021, 2020); Murgia et al. (2021); Ye and Piao (2020); Klypin et al. (2021); Hill et al. (2020); Herold et al. (2022); Herold and Ferreira (2023); Reeves et al. (2023); Jiang and Piao (2022); Simon et al. (2023); Smith et al. (2022); Nakagawa et al. (2023); Kamionkowski and Riess (2023); Niedermann and Sloth (2023); Cruz et al. (2023); Eskilt et al. (2023); Smith and Poulin (2023); Sharma et al. (2023); Efstathiou et al. (2023); Gsponer et al. (2023); Goldstein et al. (2023); Forconi et al. (2023); Fu and Wang (2024). No matter whether analyzed in light of Planck or ACT data, this model consistently yields a higher fitting value for the scalar spectral index Di Valentino et al. (2018); Ye et al. (2021, 2022); Jiang and Piao (2022); Jiang et al. (2022); Takahashi and Yin (2022); Jiang et al. (2023); Peng and Piao (2023); Forconi et al. (2023); Fu and Wang (2024). This suggests that, in EDE cosmology, resolving the Hubble tension requires moving towards the same direction favored by small-scale CMB experiments (i.e., towards larger values of $n_{s}$) [^5] Notably, evidence at a 3 standard deviation level in favor of EDE is obtained precisely from such experiments, reflecting, among other things, this preference for a higher $n_{s}$ Hill et al. (2020)..
@@ II.1 Early-time new physics
where $N_{\rm eff}$ is the effective number of relativistic particles, while $\Omega_{c}$, $\Omega_{b}$, and $\Omega_{\gamma}$ are the cold dark matter, baryons, and photons energy densities, respectively. The energy-density of various species can be accurately measured within a narrow margin of error, both through the effects in the CMB and through independent observations such as the BBN (for $\Omega_{b}$) or BAO measurements (for $\Omega_{c}$). Therefore, as also highlighted in Ref. Vagnozzi (2020), a simple (yet physically motivated) method to parameterize new physics able to increase the rate of expansion of the Universe before recombination consists of increasing the value of the effective number of relativistic particles, $N_{\text{eff}}$. Within the standard cosmological model, this parameter takes the reference value of $N_{\rm eff}=3.044$ Mangano et al. (2005); Akita and Yamaguchi (2020); Froustey et al. (2020); Bennett et al. (2021), accounting for three different families of relativistic neutrinos along with an additional contribution arising from non-instantaneous neutrino decoupling. However, given its flexibility in quantifying each relativistic component in the early Universe, this value can be changed by a wide range of phenomenological extensions spanning from modifications to the standard model of particle physics that involve new relativistic degrees of freedom active in the early Universe, to models that incorporate additional scalar degrees of freedom or significant graviton cosmic background, see e.g., Refs Giarè et al. (2021b, 2022); D’Eramo et al. (2022); Giarè et al. (2023a); Papanikolaou (2023); Garny et al. (2024).
@@ II.2 Late-time new physics
where $\Omega_{m}=\Omega_{b}+\Omega_{c}+\Omega_{\nu}$ represents the matter energy-density (encompassing baryons, cold dark matter, and neutrinos), $\Omega_{k}$ denotes the curvature density parameter, and $\Omega_{\rm{DE}}$ describes the energy-density component associated with Dark Energy (DE). Notice that I assume the DE equation of state $w\equiv P_{\rm{DE}}/\rho_{\rm{DE}}$ (i.e., the ratio between the pressure $P_{\text{DE}}$ and energy-density $\rho_{\rm{DE}}$ of DE) to be constant. Within the standard $\Lambda$CDM model of cosmology, this parameter is fixed to $w=-1$, corresponding to a cosmological constant case $\Omega_{\rm{DE}}=\Omega_{\Lambda}$. Within $\Lambda$CDM the spatial background geometry is assumed to be flat, $\Omega_{k}=0$. Therefore, if we aim to use the simple Eq. (5) to parameterize the effects of new physics beyond the standard cosmological model, we are left with only two possibilities: working on the background geometry or on the DE equation of state [^7] Notice that many other attempts to solve the Hubble tension in the context of late-time new physics have been discussed in the literature, introducing modifications to the background dynamics parameterized by Eq. (5). Just to mention a few concrete examples, a non-exhaustive list of possibilities involves considering modified gravity theories and or exotic/phenomenological theoretical frameworks such as Interacting Dark Energy or sign-switching cosmology. With no claim of covering all possibilities, for works in this direction, the interested reader can refer to, e.g., Refs. Gavela et al. (2009); Di Valentino et al. (2017); Kumar and Nunes (2017); Wang et al. (2016); Di Valentino et al. (2020a); Pan et al. (2019a); Yang et al. (2018); Murgia et al. (2016); Yang et al. (2019); Pan et al. (2019b); Di Valentino et al. (2020b); Lucca and Hooper (2020); Gómez-Valent et al. (2020); Di Valentino (2021); Yang et al. (2021a); Gariazzo et al. (2022); Bernui et al. (2023); Zhai et al. (2023); Montani et al. (2023a); Schiavone et al. (2023); Montani et al. (2023b); Akarsu et al. (2021, 2023a, 2023b); Gómez-Valent et al. (2024); Giarè et al. (2024b, c) and references therein.. Both of these parameters can be well constrained by a multitude of observational data and have been the subject of intense attention and debate for a variety of reasons [^8] See, e.g., Refs. Park and Ratra (2019); Handley (2021); Di Valentino et al. (2019); Efstathiou and Gratton (2020); Di Valentino et al. (2021c); Benisty and Staicova (2021); Vagnozzi et al. (2021a, b); Di Valentino et al. (2021d); Yang et al. (2021a); Cao et al. (2021); Dhawan et al. (2021); Dinda (2022); Gonzalez et al. (2021); Akarsu et al. (2023c); Cao and Ratra (2022); Glanville et al. (2022); Bel et al. (2022); Yang et al. (2023); Stevens et al. (2023); Favale et al. (2023) for discussions surrounding the curvature parameter and Refs. Escamilla et al. (2023); Aghanim et al. (2020b); Alam et al. (2021); Semenaite et al. (2023); Carrilho et al. (2023); Chudaykin et al. (2021); D’Amico et al. (2021); Brieden et al. (2022); Scolnic et al. (2018); Brout et al. (2022); Abbott et al. (2019); Tröster et al. (2021); Abbott et al. (2023a); Moresco et al. (2016); Vagnozzi et al. (2021b); Yang et al. (2021b); Di Valentino (2021); Moresco et al. (2022); Vagnozzi et al. (2022); Bargiacchi et al. (2022); Moresco et al. (2022); Grillo et al. (2020); Cao et al. (2022); Hogg (2023); Adame et al. (2024) for recent constraints on the DE equation of state from a multitude of astrophysical and cosmological probes.. Again, I will not take part in the debate as my goal is to parameterize new physics able to alleviate the Hubble tension and study its implications for inflation. Since one of the most typical predictions of inflation is that the observable Universe should be locally flat, I fix the background geometry to $\Omega_{k}=0$, focusing on the potential realizations of new theories capable of predicting non-standard values for $w$.